19. September 2017: Dienstag

Today’s Activities:

  • Backups:
    • Tape clones completed at 18:09 yesterday.
    • Reclaimed five (5) tapes for the vaults.
    • Disk clones completed at 08:12.
    • Tape Drive *a164c (1310270574) thinks it is loaded…
      • dump it.
      • Send dump (and email messages about not being installed) to Qualstar
      • power off, reseat, power on… seems to be working now…
      • Qualstar says dump is clean – monitor it for further issues.
    • COMIS clones completed at 12:15.
    • Backup POC cleanup – continue wipe test equipment.
  • LDAP/Accounts:
    • Working on VACC request.
    • updating carcajou caused syncrepl panic. UGH.
    • ldaprw moved to carcajou.
    • ldap servers all updated with new ACLs and Schema.

Tags:

Leave a Reply

You must be logged in to post a comment.